# Geometry-Aware Edge Pooling for Graph Neural Networks

### v1

@@ D.6 Preserving Graph Structure
Extending on the results reported in Figure 3 and Section 4.2, we further report the distribution of structure preservation measures across all pooling ratios included in our main experiment. Specifically, Figure 8 shows line plots that summarise the mean magnitude difference relative to the original graph and the normalised spectral distance between all original and pooled graphs from the NCI1 dataset in the leftmost column. The remaining plots illustrate the quantiles of the same measures split up per pooling method. Overall, these individual plots support our assessment that MagEdgePool and SpreadEdgePool consistently reach low magnitude differences and comparably low spectral distances, with these trends being more pronounced in terms of the relative difference in magnitude after pooling.
@@
Figure 8: Structure preservation for the NCI1 dataset across varying pooling ratios. Top row: The spectral distance between the normalised Laplacians of the original and the pooled graphs. Bottom row: The relative difference in magnitude, which summarises the proportional difference in structural diversity after pooling. Bold lines show the mean values of each score across graphs and thin lines the 10%, 25%, 75% and 90% quantiles.

### v2

@@ D.8 Preserving Graph Structure
Extending on the results reported in Figure˜3 and Section˜4.2, we further report the distribution of structure preservation measures across pooling ratios. Specifically, Figure˜S.8 shows line plots that summarise the mean magnitude difference relative to the original graph and the normalised spectral distance between all original and pooled graphs from the NCI1 dataset in the leftmost column. The remaining plots illustrate the quantiles of the same measures split up per pooling method. Overall, these individual plots support our assessment that MagEdgePool and SpreadEdgePool consistently reach low magnitude differences and comparably low spectral distances, with these trends being more pronounced in terms of the relative difference in magnitude after pooling. Further, we repeat the experiment reported in Figure˜3 for further datasets, specifically for DHFR, PROTEINS, and ENZYMES and summarise the results in Figure˜S.7. We observe consistent trends across these examples that support the results reported for NCI1 in Figure˜3. Specifically we find that our proposed pooling methods, MagEdgePool and SPreadEdgePool, reach the lowest magnitude differences across pooling ratios and datasets, and that this corresponds to low spectral distances between the pooled and original graphs.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
@@
Figure S.7: Structure preservation for all graphs in the ENZYMES, DHFR, and PROTEINS datasets across pooling ratios. Left: The spectral distance between the normalised Laplacians of the original and the pooled graphs. Right: The relative difference in magnitude, summarising proportional differences in structural diversity after pooling. Violin plots show the variability across graphs at pooling ratio 0.5.
Figure S.8: Structure preservation for the NCI1 dataset across pooling ratios. Top row: The spectral distance between the original and pooled graphs. Bottom row: The relative difference in magnitude. Bold lines show the mean values of each score across graphs and thin lines the 10%, 25%, 75% and 90% quantiles.